# A Lean Dataset for International Math Olympiad: Small Steps towards Writing Math Proofs for Hard Problems

### v1

@@ Abstract
Using AI to write formal proofs for mathematical problems is a challenging task that has seen some advancements in recent years. Automated systems such as Lean can verify the correctness of proofs written in formal language, yet writing the proofs in formal language can be challenging for humans as well as automated systems. The miniF2F dataset, the community’s benchmark for automated theorem proving, has 20 IMO problems in its testing set, yet formal proofs are available only for 7 of these problems (3 of which are written by mathematicians). The model with best accuracy can only prove 4 of these 20 IMO problems, from 1950s and 60s, while its training set is a secret. In this work, we write complete, original formal proofs for the remaining 13 IMO problems in Lean along with 3 extra problems from IMO 2022 and IMO 2023. This effort vastly expands the availability of proof currently in the public domain by creating 5,150 lines of lean proof. The goal of the paper is to pave the way for developing AI models that can automatically write the formal proofs for all the IMO problems in miniF2F and beyond. In this pursuit, we devise a method to decompose the proof of these problems into their building blocks, constructing a dataset of about 900 lemmas with 25,500 lines of Lean code. These lemmas are not trivial, yet they are approachable, providing the opportunity to evaluate and diagnose the failures and successes of AI models. We then evaluate the ability of GPT-4 in writing formal proofs for these lemmas. This involves zero shot prompting, Chain-of-Thought reasoning with expert feedback and lemma retrieval. In evaluating the responses, we also analyze the confounding factor of LLM’s ability to write the proofs in natural language vs Lean language. From all these evaluations, two clusters of IMO problems emerge: one group are the problems prior to 1970s where LLMs have a higher success rate in writing a proof for their decomposed lemmas, they can explain a correct proof in natural language, and when they fail, expert feedback and lemma retrieval improves their answers. The second group of problems are the IMO problems from the more recent years, their proofs are longer, subjects are more challenging, the success rate of LLM in proving the decomposed lemmas is small, LLM cannot explain a correct proof for them in natural language, and finally, lemma retrieval and CoT rarely have a positive impact on the accuracy of their formal proofs.
@@ 1 Introduction
Machine learning methods have made significant progress in many different domains. This progress can often be described as the automation of learning from data. Availability of high quality data is crucial for the community to develop better models and methods. ImageNet [10], for example, played a significant role in the development of ML methods in computer vision. More recent models such as GPT and Llama rely on training set purportedly almost as large as the entire contents of the Web. On the other hand, the failures of these models on topics such as math are often explained via lack of high-quality materials on those topics in their training set. This issue is partially addressed by curation of relevant data, e.g., [17] recruits human labelers and AlphaGeometry [29] generates synthetic data. For the topic of automated theorem proving in formal language, high quality data is so scarce, to the extent that the formal proofs are not available for almost half of the problems in the miniF2F dataset, the benchmark of the community. So, when automated methods fail to write a correct proof for the problems in the benchmark dataset, it is not clear where they have failed, and what the correct answer looks like. In this work, we study 16 IMO problems, in number theory and algebra. 13 of the 16 problems are from the miniF2F dataset and 12 of the 16 do not have a formal proof available in the public domain. Moreover, for all the 40 IMO problems in the miniF2F dataset, we contribute the complete, original, formal proofs in Lean language.
@@
Solving mathematical problems is a challenging task for machine learning. There are many motivations for doing math with AI as we will discuss further in Appendix B. One motivation is the continuation of creating machines that can outperform humans in a given task, such as in the game of chess [19] and two decades later in the game of Go [25]. Currently, there is a 10 million dollar prize for an AI system that can participate in the International Math Olympiad and win a gold medal: the AIMO Prize [1]. There is also the IMO Grand Challenge [2] which specifically requires the model to write its proof in Lean language. Lean is a functional programming language and a proof assistant [9] used mostly by mathematicians to write and verify mathematical proofs. The Lean system makes it possible to verify the correctness of a proof instantly. This has given rise to a field known as formal mathematics. Lean is also used by Amazon AWS for formal verification of software [8].
1.1 IMO: an unknown testing set that gets harder every year
@@
These conditions, in the language of machine learning, implies being evaluated on an unknown testing set with samples that are more challenging than the samples in the training set. In the machine learning literature, however, having standard benchmarks with fixed and accessible testing sets is essential for a large community to be able to work towards common goals. This approach of using a fixed testing set has worked reasonably well in some ML fields [22], mostly because the training sets of the models are fixed, too, and in creating these datasets, specific procedures are followed to ensure samples from training sets do not appear in the testing sets. However, with the rise of large language models, it has become increasingly difficult to evaluate how much of a model’s success is merely direct retrieval of the correct answer from its training set. Although for solving mathematical problems, the miniF2F [34] dataset is the common benchmark with a fixed testing set, the training set of LLMs is sometimes unknown; when it is known, it includes all the relevant content available on the Internet, (e.g., all the available code on Github, arXiv, Stack Exchange, and elsewhere), which is an enormous corpus, hard to search through and compare against the testing sets. When evaluating LLMs developed on such training sets, we often use a fixed testing set. But, IMO evaluates on an unknown testing set that gets harder every year.
To prepare an AI model that can do well in IMO, we need to rely on rigorous evaluation techniques that can tell us whether the model can generalize well on an unknown testing set. In some applications, the distinction between generalization and retrieval might be besides the point. As long as a model performs the task correctly, one might not be too keen to evaluate what percentage of its correct answers is a direct retrieval from its training set. However, when we are developing a model that is intended to be evaluated on an unknown testing set, we will need to rely on clearer evaluation methods.
@@ 1.2 Where we are
Let us now take a closer look at the current benchmark: miniF2F dataset. This dataset consists of mathematical theorems with varying degrees of difficulty. Out of the 244 theorems in its testing set, 20 of them are IMO problems, and the difficulty levels of those IMO problems also vary since they are taken from years 1959 to 2019. In terms of complexity, some of the theorems in the dataset can be proved with one line of formal proof while others may need more than 500 lines of formal proof (see Table 1). Success in solving a high percentage of the less complex problems is not the same as success in solving the same percentage of the more complex problems. Hence, merely reporting the accuracy by the percentage of the testing set is not insightful enough.
LEGO-Prover [30], the current state of the art method on miniF2F, is able to prove only one of the IMO problems in this dataset: the one from 1959, which is the first problem of the first IMO problem implying that it is the easiest of all IMO problems proved in less than 10 lines of code. Other ML methods have proved more of the IMO problems in the miniF2F, e.g., the HTPS method by [16] claims proving 10 IMO problems (not all from the miniF2F), but those problems are either from 1960s or the ones for more recent years have a simplified or erroneous problem statement. The method by [21] also proves two adapted (simplified) IMO problems.
@@ 1.3 Our approach
We provide the formal proofs for these problems and further evaluate the ability of GPT-4 in writing these proofs. We decompose each of the formal proof steps into small lemmas and report that GPT-4 cannot write a correct formal proof for almost all of the lemmas. Writing the formal proof for these problems can be viewed from the perspective of planning and execution: one would need to prove a certain chain of lemmas (execution), and put those lemmas together (planning) in a coherent way to compose the final proof. The planning aspect of writing formal proofs is challenging, and in the literature, often the plan, i.e., the proof sketch or the approach, is taken from the informal proofs written by humans [14].
We diagnose the failures of GPT-4 by evaluating and labeling its proofs manually. We further guide the model by pointing out its mistakes. Giving feedback to the model is mostly unfruitful, especially on the topics that GPT-4 seems to be less educated. Finally, we perform source criticism by comparing the proofs written by GPT-4 against the contents of the Web, as a proxy for its training set. This systematic study provides insights as to what fraction of the answers of GPT-4 can be considered direct retrieval from its training set. We report a strong correlation between the correctness of the answers of GPT-4 and the existence of proofs on the Web.
@@ 1.4 Our contributions
    We provide the formal proofs for 12 IMO problems that do not have a formal proof in Lean. Nine of these problems are from the miniF2F dataset. The other three are from IMO 2022 and 2023. We specifically chose the problems that do not have a formal Lean proof in the public domain. Our proofs are written in Lean3 and 4 and consist of 3,100 lines of code.
@@
    We decompose the proofs for 12 problems into small lemmas, creating a dataset of more than 900 lemmas that are still challenging for AI models, yet they are easier and more approachable. Our method can be extended to all the IMO problems that we have formalized.
@@
    We evaluate the ability of GPT-4 in writing the formal proofs for the lemmas in on our dataset. We perform a thorough human review of the GPT-4’s informal proofs as well as its formal proofs. For incorrect proofs, we consider Chain-of-Thoughts techniques to guide the GPT-4. We further quantify the human effort required to repair the incorrect proofs.
@@ 2 Dataset of approachable lemmas: Stepping stone towards solving challenging IMO problems
Table 1 describes the 12 IMO problems that we study in this paper. The concepts used in these problems are described in further detail in Table C1. In appendices D and F, we provide the datasheet and license information about our dataset.
We report that GPT-4 cannot write a correct formal proof for any of these problems. In natural language, though, it can describe a sensible approach for proving the first 6 IMO problems. For the IMO problems between 1984 and 2023, its approach in natural language is also wrong.
@@
# Year Problem Topic in miniF2F Lean proof available # of lemmas # of lines of Lean4 code
1 1959 p1 number theory Yes No 3 9
2 1960 p2 algebra Yes No 8 40
@@
6 1983 p6 algebra Yes No 55 180
7 1984 p6 number theory Yes No 68 380
@@
total 907 3,149
Making these formal proofs accessible to the research community can be helpful in different ways. First, every one, including non-mathematicians, will know the work it takes to write a formal proof for these problems. Second, these formal proofs can also be used to identify possible mistakes or shortcomings in the informal proofs that might be circulating. For example, [5] reports that GPT-4 can write a correct informal proof for IMO 2022 P2. However, when we write the formal proof for the same problem, it becomes obvious that the informal proof was incorrect and missing a crucial step. Third, some of the formal proofs in the literature simplify the original IMO problem by altering the problem statement, but we remain faithful to the original problem statements of IMO. For example, [31] showcases the proof for two shortlisted IMO problems. In the first problem, IMO 2009 P3, there is a mistake in the problem statement, and the proof merely uses the error to prove False. For the second problem, IMO 2016 P5, the statement is altered from proving “there exists infinitely many solutions" to “there exists a solution".
@@ 2.1 Decomposition of the proofs into lemmas
In the next step, we decompose the proofs for each of these problems into small lemmas. For example, the proof might involve showing that given a set of constraints, natural number $p$ cannot be larger than 4. Then, use another series of deductions to show that $p$ cannot be smaller than 2. And then use these two results to conclude that $p$ must be either $2$ or $3$. Appendix E provides a more detailed discussion about decomposition of lemmas and some upper bounds on how many proofs can be extracted from $n$ lines of proof. Moreover, Table E2 provides some statistics about the length of the proofs for the lemmas in our dataset.
All of the 907 lemmas in our dataset have a formal proof in Lean. Some of the easier lemmas are as the following:
@@
  • $q$ and $r$ are natural numbers and their product is 5, prove that either $q$ or $r$ must be 1.
@@
The lemmas that are harder to solve are also broken into smaller pieces, so the dataset includes both the hard version of such lemmas plus the smaller lemmas needed to prove them. Later in section 4, we will go through a case study of such lemma to give a better understanding of the dataset and what it takes to prove the lemmas.
2.2 Excluding the problems that native Lean methods can prove
We intentionally exclude lemmas that can be solved automatically via Lean’s native methods such as hint, linarith, exact?, simp, omega, ring, norm_cast and norm_num. This is because we consider those solvers efficient and we do not find it essential to develop automated methods doing the same job as those solvers. We have empirically seen that GPT-4 fails to write a correct formal proof for many of those easy problems as well. Nevertheless, our main objective here is not to evaluate GPT-4, but to create a dataset that can be used as a stepping stone for automated systems. If a task can already be performed adequately and fast by some open source solver, our view is that it would be better to use such a solver and even incorporate it into a more sophisticated system.
3 GPT-4’s ability to solve the problems
3.1 Prompting GPT-4

### v2

@@ Abstract
Using AI to write formal proofs for mathematical problems is a challenging task that has seen some advancements in recent years. Automated systems such as Lean can verify the correctness of proofs written in formal language, yet writing the proofs in formal language can be challenging for humans and machines. The miniF2F benchmark has 20 IMO problems in its test set, yet formal proofs are available only for 6 of these problems (3 of which are only written by mathematicians). The model with best accuracy can only prove 2 of these 20 IMO problems, from 1950s and 60s, while its training set is a secret. In this work, we write complete, original formal proofs for the remaining IMO problems in Lean along with 3 extra problems from IMO 2022 and 2023. This effort expands the availability of proof currently in the public domain by creating 5,880 lines of Lean proof. The goal of the paper is to pave the way for developing AI models that can automatically write the formal proofs for all the IMO problems in miniF2F and beyond by providing an evaluation benchmark. In this pursuit, we devise a method to decompose the proofs of these problems into their building blocks, constructing a dataset of 1,329 lemmas with more than 40k lines of Lean code. These lemmas are not trivial, yet they are approachable, providing the opportunity to evaluate and diagnose the failures and successes of AI models. We evaluate the ability of the SOTA LLMs on our dataset and analyze their success and failure modes from different perspectives. Our dataset and code is available at: https://github.com/roozbeh-yz/IMO-Steps.
@@ 1 Introduction
Machine learning methods have made significant progress in many different domains. This progress can often be described as the automation of learning from data. The availability of high quality data is crucial for the community to develop better models and methods. ImageNet (Deng et al., 2009), for example, played a significant role in the development of ML methods in computer vision. More recent models such as OpenAI’s o1 and Llama rely on training sets purportedly almost as large as the entire contents of the Web. On the other hand, the failures of these models on topics such as math are often explained via the lack of high-quality materials on those topics in their training set. This issue is partially addressed by curation of relevant data, e.g., Lightman et al. (2023) recruits human labelers and AlphaGeometry (Trinh et al., 2024) generates synthetic data. For the topic of automated theorem proving in formal language, high-quality data is so scarce, to the extent that the formal proofs are not available for almost half of the problems in the miniF2F dataset, the benchmark of the community. So, when automated methods fail to write a correct proof for the problems in the benchmark dataset, it is not clear where they fail, and what the correct answer looks like. In this work, we contribute the complete, original, formal proofs in Lean language for all the 14 IMO problems in the test set of miniF2F that do not have a formal proof publicly available. Moreover, we study 13 IMO problems in detail, breaking them down into their building blocks, leading to a dataset of 1,330 lemmas. We then evaluate the ability of SOTA LLMs on our dataset. The results reveal the abilities and inabilities of these models in proving the building blocks of IMO problems.
@@
Solving mathematical problems is a challenging task for machine learning. There are many motivations for doing math with AI as we will discuss further in the Appendix J. One motivation is to continue creating machines that can outperform humans in a given task, such as in the game of chess (Pandolfini, 1997) and two decades later in the game of Go (Silver et al., 2017). Currently, there is a 10-million dollar prize for an AI system that can participate in the International Math Olympiad and win a gold medal: the AIMO Prize. There is also the IMO Grand Challenge which specifically requires the model to write its proof in Lean language. Lean is a functional programming language and a proof assistant (de Moura et al., 2015) used mostly by mathematicians to write and verify mathematical proofs. The Lean system makes it possible to verify the correctness of a proof almost instantly. This has given rise to a field known as formal mathematics. Lean is also used by Amazon AWS for formal verification of software (Cutler et al., 2024).
Our approach in this paper is to provide a stepping stone for proving IMO problems in the miniF2F dataset and beyond. With the help of our dataset and the complete proofs that we contribute to the test set of miniF2F benchmark, we hope to enable researchers to identify and better understand the weaknesses of their models. Consider the case of a math teacher who encounters a student who cannot solve a hard problem. The teacher might ask the student to solve easier problems first, such as the building blocks of that hard problem. How much the student can solve the easier problems will help the teacher evaluate the student’s strength and weaknesses. Likewise, the building blocks we provide in our dataset can be considered stepping stones for evaluation purposes. For further illustration, we can also break down the evaluation into scenarios such as the following: 1- The student cannot provide a correct answer for any of the easy problems. The teacher might start asking even simpler questions. 2- The student can provide a correct answer for some of the easy problems. The teacher may be able to identify the underlying weaknesses of the student. Perhaps the student has not been exposed to certain subjects. Or, maybe the student has not learned certain reasoning methods. 3- The student can solve all of the easy problems. Then the teacher might try to teach the student how to put the answers for the easy problems together to compose the answer for the more complex problems.
The lemmas that we provide in our dataset are analogous to the easy problems that the teacher uses in the example above. Therefore, our lemmas are designed to better evaluate the models and provide a way to diagnose their failures. We do not suggest that people should fine-tune their models on the complete proofs that we provide. Rather, we suggest creating a clear separation between the training set of the models and the test set of miniF2F. By making these proofs public, we anticipate the community will move towards better evaluation practices.
1.1 IMO: an unknown test set that gets harder every year
@@
These conditions, in the language of machine learning, implies being evaluated on an unknown test set with samples that are more challenging than the samples in the training set. In the machine learning literature, however, having standard benchmarks with fixed and accessible test sets is essential for a large community to be able to work towards common goals. This approach of using a fixed test set has worked reasonably well in some ML fields (Recht et al., 2019), mostly because the training sets of the models are fixed, too, and in creating these datasets, specific procedures are followed to ensure samples from training sets do not appear in the test sets. However, with the rise of large language models, it has become increasingly difficult to evaluate how much of a model’s success is merely direct retrieval of the correct answer from its training set. Although for solving mathematical problems, the miniF2F (Zheng et al., 2021) dataset is the common benchmark with a fixed test set, the training set of LLMs is sometimes unknown; when it is known, it includes all the relevant content available on the Internet, (e.g., all the available code on Github, arXiv, Stack Exchange, and elsewhere), which is an enormous corpus, hard to search through and compare against the test sets. When evaluating LLMs developed on such training sets, we often use a fixed test set. But, IMO evaluates on an unknown test set that gets harder every year.
To prepare an AI model that can do well in IMO, we need to rely on rigorous evaluation techniques that can tell us whether the model can generalize well on an unknown test set. In some applications, the distinction between generalization and retrieval might be besides the point. As long as a model performs the task correctly, one might not be too keen to evaluate what percentage of its correct answers is a direct retrieval from its training set. However, when we are developing a model that is intended to be evaluated on an unknown test set, we will need to rely on clearer evaluation methods.
@@ 1.2 Where we are
Let us now take a closer look at the current benchmark: miniF2F dataset. This dataset consists of mathematical theorems with varying degrees of difficulty. Out of the 244 theorems in its test set, 20 of them are IMO problems, and the difficulty levels of those IMO problems also vary since they are taken from years 1959 to 2019. In terms of complexity, some of the theorems in the dataset can be proved with one line of formal proof while others may need more than a thousand lines of formal proof (see Table 1). Success in solving a high percentage of the less complex problems is not the same as success in solving the same percentage of the more complex problems. Hence, merely reporting the accuracy by the percentage of the test set is not insightful enough.
LEGO-Prover (Wang et al., 2024), the current peer-reviewed state of the art method on miniF2F, is able to prove only one of the IMO problems in this dataset: the one from 1959, which is the first problem of the first IMO competition implying that it is the easiest of all IMO problems proved in less than 10 lines of code. Other ML methods have proved more of the IMO problems in the miniF2F, e.g., the HTPS method by Lample et al. (2022) claims proving 10 IMO problems (not all from the miniF2F), but those problems are either from 1960s or the ones for more recent years have a simplified or erroneous problem statement. The method by Polu et al. (2022) also proves two adapted (simplified) IMO problems.
@@ 1.3 Our approach
We provide the formal proofs for these problems and further evaluate the ability of o3-mini (OpenAI, 2025) in writing these proofs. We decompose each of the formal proof steps into small lemmas and report that o3-mini cannot write a correct formal proof for almost all of the lemmas. Writing the formal proof for these problems can be viewed from the perspective of planning and execution: one would need to prove a certain chain of lemmas (execution), and put those lemmas together (planning) in a coherent way to compose the final proof. The planning aspect of writing formal proofs is challenging, and in the literature, often the plan, i.e., the proof sketch or the approach, is taken from the informal proofs written by humans (Jiang et al., 2022).
We diagnose the failures of o3-mini by evaluating and labeling its proofs manually. We further guide the model by pointing out its mistakes. Giving feedback to the model is mostly unfruitful, especially on the topics that o3-mini seems to be less educated.
@@ 1.4 Our contributions
    We provide the formal proofs for 17 IMO problems that do not have a formal proof in Lean. 14 of these problems are from the miniF2F dataset. The other three are from IMO 2022 and 2023. We specifically chose the problems that do not have a formal Lean proof in the public domain. Our proofs are written in Lean 4 and consist of 5,790 lines of code.
@@
    We decompose the proofs of 13 IMO problems into small lemmas, creating a dataset of 1,329 lemmas that are still challenging for AI models, yet they are easier and more approachable.
@@
    We evaluate the ability of the stat-of-the-art LLMs (Goedel-Prover (Lin et al., 2025), DeepSeek-Prover-V1.5-RL (Xin et al., 2024b), and OpenAi’s o3-mini) in writing the formal proofs for the lemmas in on our dataset. We further perform a human review of o3-mini’s informal and formal proofs after 10 rounds of expert feedback. We categorize the issues in the incorrect proofs providing deep insights about what goes wrong. We also analyze the correct proofs observing that they are limited by length.
@@ 2 Dataset of approachable lemmas: Stepping stone towards solving challenging IMO problems
Table 1 describes the 13 IMO problems that we study in this paper. The concepts used in these problems are described in further detail in Table B1. In appendices I and H, we provide the datasheet and license information about our dataset.
Note that we provide the Lean proofs for all the IMO problems in the test set of miniF2F, but we limit our study to a subset of these problems listed in Table 1. Table C2 provides information about the proof of all the problems in our dataset including the ones that we do not perform lemma decomposition. Additionally, Table D3 provides information about the lengths of the proofs for the lemmas in our proposed dataset.
@@
# Year Problem Topic in miniF2F Lean proof publicly available # of lemmas # of lines of Lean4 proof
1 1959 p1 number theory Yes Yes 4 9
2 1960 p2 algebra Yes Yes 9 40
@@
6 1983 p6 algebra Yes No 53 180
7 1984 p6 number theory Yes No 64 380
8 1985 p6 number theory Yes No 427 1,310
@@
total 1,329 4,459
Making these formal proofs accessible to the research community can be helpful in different ways. First, every one, including non-mathematicians, will know the work it takes to write a formal proof for these problems. Second, these formal proofs can be used to identify possible mistakes or shortcomings in the informal proofs that might be circulating. For example, Bubeck et al. (2023) reports that GPT-4 can write a correct informal proof for IMO 2022 P2. However, when we write the formal proof for the same problem, it becomes obvious that the informal proof was missing a crucial step. Third, some of the formal proofs in the literature simplify the original IMO problem by altering the problem statement, but we remain faithful to the original problem statements of IMO. For example, Xin et al. (2024a) showcases the proof for two shortlisted IMO problems. In the first problem, IMO 2009 P3, they have a mistake in the problem statement, and the proof merely uses the error to prove False. For the second problem, IMO 2016 P5, the statement is altered from proving “there exists infinitely many solutions” to “there exists a solution”.
@@ 2.1 Decomposition of the proofs into lemmas
In the next step, we decompose the proofs for each of these problems into smaller lemmas. For example, the proof might involve showing that given a set of constraints, natural number $p$ cannot be larger than 3. Then, use another series of deductions to show that $p$ cannot be smaller than 2. And then use these two results to conclude that $p$ must be either $2$ or $3$. Section 4 provides a more detailed discussion about decomposition of lemmas and an upper bound on how many proofs can be extracted from $n$ lines of proof. Moreover, Table D3 provides some statistics about the length of the proofs for the lemmas in our dataset.
All of the 1,329 lemmas in our dataset have a formal proof in Lean 4, version 4.17. Some of the easier lemmas are as the following:
@@
  • $f$ is a real-valued function that is strictly monotonic and upper bounded by $a$, prove that $a$ cannot be in the range of $f$.
@@
The lemmas that are harder to solve are also broken into smaller pieces, so the dataset includes both the hard version of such lemmas plus the smaller lemmas needed to prove them.
3 Evaluating the ability of LLMs in proving our lemmas
We first evaluate o3-mini in great depth, then we analyze the latest SOTA models on our dataset. Finally, we analyze the success and failure cases of the LLMs.
3.1 o3-mini’s ability in formal and natural language